OpenStudy (anonymous):

What is the simplified form of the following expression? sqrt of 1/121 A.2/121 B.1/242 C.11 D.1/11

Nnesha (nnesha):

\[\huge\rm \sqrt{\frac{ 1 }{ 121 }}\] like this ?

OpenStudy (anonymous):

Thats just like it :)

Nnesha (nnesha):

remember \[\huge\rm \sqrt{\frac{ a }{ b }}= \frac{ \sqrt{a} }{ \sqrt{b} }\] take square root of the numerator and denominator

OpenStudy (anonymous):

That means......1/11 ?

Nnesha (nnesha):

:=)

OpenStudy (anonymous):

Thanks! :)

Nnesha (nnesha):

my pleasure :=)